Pablo is putting 10 books in a row on his bookshelf. He will put one of the books, Pride and Predjudice, in the first spot. He will put another of the books, Little Women, in the last spot. In how many ways can he put the books on the shelf?
Answer by ikleyn(52776)

.
Pablo is putting 10 books in a row on his bookshelf. He will put one of the books, Pride and Predjudice, in the first spot.
He will put another of the books, Little Women, in the last spot. In how many ways can he put the books on the shelf?
~~~~~~~~~~~~~~~~~~~~~~~

The context (the hidden part of the problem) is that the rest of 10 books, 
i.e. 10-2 = 8 books, can be located in any of 8 remaining spots.


8 books can be placed in 8 spots by 8! = 8*7*6*5*4*3*2*1 = 40320 different ways (permutations).    ANSWER

Solved, answered and explained.